OpenStudy (anonymous):

Richard's estate is to be split equally among his three children. If his estate is worth $1 3/4 million, how much will each child inherit?

OpenStudy (anonymous):

1.75 / 3 = 0.5833333 million or $583,333.33

OpenStudy (anonymous):

\[\frac{1.75 \times 10^{6}}{3} = 5.833 \times 10^{5} = $583,333\]